OpenStudy (anonymous):

priya is raising money for a charity by selling tomato plants that she has grown from seeds. She sells 48 tomato plants for £1.35 each. Priya keep 15 percent of the money she gets to pay for the growbags and seeds that she used. How much money did she send to the charity? HELP!

OpenStudy (anonymous):

Somebody help me

OpenStudy (ash2326):

Priya sells 48 tomato plants for 1.35 each. her earnings= \(48\times 1.35\)=\(64.8\) She keeps 15% percent of the money. So she sends 85% of the money to charity \[= 85\%64.8=64.8\times \frac{85}{100}= 55.08 \]

OpenStudy (ash2326):

loteviex you got it?

OpenStudy (anonymous):

yes thanks

OpenStudy (ash2326):

welcome:)